OpenStudy (anonymous):

Solve Y/-2 − 3 = 12 30 −30 18 −18

OpenStudy (anonymous):

y/-2-3=12 y/-2=15 y=-30

OpenStudy (anonymous):

-30

OpenStudy (anonymous):

y/-2=12+3 y=15*-2 y=-30